OpenStudy (anonymous):

Choose the correct GCF for the following binomial: 6a^4b^2 + 9a^3b

sam (.sam.):

6a^(4)b^(2)+9a^(3)b Factor out the GCF of 3a^(3)b from 6a^(4)b^(2)+9a^(3)b. 3a^(3)b(2ab+3)

OpenStudy (anonymous):

whats the answer?

OpenStudy (mertsj):

Sam gave you the answer. How is it you ask a question, Sam answers it, and you can't identify the answer?

OpenStudy (anonymous):

cause thats not how long the answer would be i have a guess at it... I think its 3a^3b

OpenStudy (mertsj):

What does GCF mean?

OpenStudy (anonymous):

greatest common factor

OpenStudy (mertsj):

What does that mean?

OpenStudy (mertsj):

Is 10 the GCF of 2 and 5?

OpenStudy (anonymous):

No

OpenStudy (mertsj):

What is the GCF of 2 and 5?

OpenStudy (anonymous):

5

OpenStudy (mertsj):

5 does not divide into 2

OpenStudy (anonymous):

actually i dont think there is 1

OpenStudy (anonymous):

unless its 1?

OpenStudy (mertsj):

Every number has 1 as a factor so the GCF of 2 and 5 is 1.

OpenStudy (mertsj):

What is the GCF of 6 and 9?

OpenStudy (anonymous):

3

OpenStudy (mertsj):

What is tHE GCF of a^4 and a^3?

OpenStudy (anonymous):

a^1

OpenStudy (mertsj):

Will a^2 divide into a^4 and a^3?

OpenStudy (anonymous):

into 4 yes but not into 3

OpenStudy (mertsj):

Are you remembering that when you divide exponentials, you subtract the exponents?

OpenStudy (anonymous):

nope

OpenStudy (mertsj):

\[\frac{a^4}{a^3}=a^{4-3}=a^1\]

OpenStudy (mertsj):

Do you know that?

OpenStudy (anonymous):

no i didnt know that

Directrix (directrix):

Didn't .Sam answer the question? GCF is --> 3a³b

OpenStudy (anonymous):

no he didnt